# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ 1 Introduction
The LASSO (Tibshirani, 1996) is perhaps the most commonly employed approach to handle regressions with a large number of predictors. From a theoretical standpoint, its effectiveness in terms of estimation, prediction, and feature selection is contingent upon either orthogonality or reasonably weak correlation among predictors (see Zhao and Yu, 2006; Bickel et al., 2009; Negahban et al., 2012; Hastie, 2015). This hinders the use of the LASSO for the analysis of economic time series data, which are notoriously characterized by intrinsic multicollinearity; that is, by predictor correlations at the population level (Forni et al., 2000; Stock and Watson, 2002a; De Mol et al., 2008; Medeiros and F.Mendes, 2012). A common procedure to address this issue is to model multicollinearity and remove it, as proposed, e.g., by  Fan et al. (2020), who filter time series using common factors and then apply the LASSO to the filtered residuals. However, mitigating or even eliminating multicollinearity is not the end of the story, as effectiveness of the LASSO can also be affected by spurious correlations. These occur when predictors are orthogonal or weakly correlated at the population level, but a lack of sufficient independent replication (lack of degrees of freedom) introduces correlations at the sample level, potentially leading to false scientific discoveries and incorrect statistical inferences (Fan and Zhou, 2016; Fan et al., 2018). This issue has been broadly explored in ultra-high dimensional settings, where the number of predictors can vastly exceed the available sample size (Fan et al., 2014). We argue that in time series data, a shortage of independent replication can be due not only to a shortage of available observations but also to serial correlation.
@@
Second, we propose an approach that, using a filter similar to that proposed by Fan et al. (2020), rescues the performance of the LASSO in the presence of serially correlated predictors. Our approach, which we name ARMAr-LASSO (ARMA residuals LASSO), relies upon a working model where, instead of the observed predictor time series, we use as regressors the residuals of ARMA processes fitted on such series, augmented with lags of the dependent variable. We motivate our choice of working model and provide some asymptotic arguments concerning limiting distribution and feature selection consistency. Next, we employ the mixingale and near-epoch dependence framework (Davidson, 1994; Adamek et al., 2023) to prove oracle inequalities for the estimation and forecast error bounds of the ARMAr-LASSO, while simultaneously addressing the issue of estimating ARMA residuals. To complete the analysis, we use simulations to validate and generalize theoretical results. Furthermore, we apply our methodology to a high-dimensional dataset for forecasting the consumer price index in the Euro Area. Simulations and empirical exercises demonstrate that the ARMAr-LASSO produces more parsimonious models, better coefficient estimates, and more accurate forecasts than LASSO-based benchmarks. Notably, both theoretical and numerical results concerning our approach hold even in the presence of factor-induced multicollinearity, provided that the idiosyncratic components are orthogonal or weakly correlated processes exhibiting serial correlation.
Our work complements the vast literature on error bounds for LASSO-based methods in time series analysis, which addresses estimation and forecast consistency in scenarios with autocorrelated errors and autoregressive processes (Bartlett, 1935; Granger and Newbold, 1974; Granger et al., 2001; Wang et al., 2007; Nardi and Rinaldo, 2011; Uematsu and Tanaka, 2019; Babii et al., 2022; Chronopoulos et al., 2023; Baillie et al., 2024). Such scenarios are ubiquitous, e.g., they are easily found in US and Euro Area monthly macroeconomic data (see McCracken and Ng, 2016 and Proietti and Giovannelli, 2021). Moreover, our methodology aligns with the prevailing literature on pre-whitening filters, which are commonly employed to mitigate autocorrelation in the error component or to enhance cross-correlation analysis. For instance, Robinson (1988) suggests obtaining residuals from partial linear models and then applying least squares to estimate a non-parametric component. Belloni et al. (2013) propose a similar approach in high-dimensional settings, using LASSO on the residuals from a parametric component. Hansen and Liao (2019) show that, in panel data model inference, applying a PCA filter to eliminate multicollinearity and using LASSO on its residuals can produce reliable confidence intervals. Finally, as mentioned, Fan et al. (2020) illustrate how latent factors create strong dependencies among regressors, complicating feature selection, while selection improves on idiosyncratic components.
@@
The remainder of the article is organized as follows. Section 2 introduces the problem setup and our results concerning the link between serial correlations and spurious correlations. Section 3 introduces the ARMAr-LASSO and explores its theoretical properties. Section 4 presents simulations and real data analyses to evaluate the performance of our proposed methodology. Section 5 provides some final remarks.
We summarize here some notation that will be used throughout. Bold letters denote vectors, for example $\mathbf{a}=(a_{1},\dots,a_{p})^{\prime}$. Supp$(\mathbf{a})$ denotes the support of a vector, that is, $\{i\in\{1,\dots,p\}:a_{i}\neq 0\}$, and $|\text{Supp}(\mathbf{a})|$ the support cardinality. The $\ell_{q}$ norm of a vector is $||\mathbf{a}||_{q}\coloneqq\mathopen{}\mathclose{{}\left(\sum_{j=1}^{p}|a_{j}|^{q}}\right)^{1/q}$ for $0<q<\infty$, with $||\mathbf{a}||_{q}^{k}\coloneqq\mathopen{}\mathclose{{}\left(\sum_{j=1}^{p}|a_{j}|^{q}}\right)^{k/q}$, and with the usual extension $||\mathbf{a}||_{0}\coloneqq|\text{Supp}(\mathbf{a})|$. Bold capital letters denote matrices, for example $\mathbf{A}$, where $\mathopen{}\mathclose{{}\left(\mathbf{A}}\right)_{ij}=a_{ij}$ is the $i$-row $j$-column element. Furthermore, $\boldsymbol{0}_{p}$ denotes a $p$-length vector of zeros, $\boldsymbol{I}_{p}$ the $p\times p$ identity matrix, and $\text{Sign}(r)$ the sign of a real number $r$. $\mathopen{}\mathclose{{}\left\lceil x}\right\rfloor$ indicates that $x$ has been rounded to the nearest integer. To simplify the presentation, we frequently use $C$ to indicate arbitrary positive finite constants.
@@ 2 Problem Setup
where $\mathbf{x}_{t}=(x_{1,t},\ldots,x_{n,t})^{\prime}$ is a $n\times 1$ vector of predictors, $\boldsymbol{\alpha}^{*}$ is a $n\times 1$ unknown $s$-sparse vector of regression coefficients, i.e. $||\boldsymbol{\alpha}^{*}||_{0}=s<n$, and $\varepsilon_{t}$ is an error term. We impose the following assumptions on the processes $\mathopen{}\mathclose{{}\left\{\mathbf{x}_{t}}\right\}$ and $\mathopen{}\mathclose{{}\left\{\varepsilon_{t}}\right\}$.
@@
(b) $u_{i,t}\perp u_{j,t-l}$ for any $i\neq j,\ t$ and $l\neq 0$; and $u_{i,t-l}\perp\omega_{t}$ for any $i,t$ and $l$.
The innovation processes $u_{i,t}\sim w.n.(0,\sigma_{i}^{2})$, $\omega_{t}\sim w.n.(0,\sigma_{\omega}^{2})$ are sequences of zero-mean white noise ($w.n.$) with finite variances (see  Reinsel, 1993, ch. 1.2 for details).
@@
There are several approaches to estimate a sparse $\boldsymbol{\alpha}^{*}$ (Zhang and Zhang, 2012; James et al., 2013); here we focus on the LASSO estimator (Tibshirani, 1996) given by $\widehat{\boldsymbol{\alpha}}=\underset{\boldsymbol{\alpha}\in\mathbb{R}^{n}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{}\left\{\frac{1}{2T}||\mathbf{y}-\mathbf{X}^{\prime}\boldsymbol{\alpha}||_{2}^{2}+\lambda||\boldsymbol{\alpha}||_{1}}\right\},$ where $\mathbf{y}=(y_{1},\dots,y_{T})^{\prime}$ is the $T\times 1$ response vector, $\mathbf{X}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$ is the $n\times T$ design matrix, and $\lambda>0$ is the weight of the $\ell_{1}$ penalty and must be “tuned” to guarantee that regression coefficient estimates are effectively shrunk to zero – thus ensuring predictor, or feature, selection.
@@
However, linear associations among predictors are well known to affect LASSO performance. Bickel et al. (2009); Bühlmann and van de Geer (2011) and Negahban et al. (2012) have shown that the LASSO estimation and prediction accuracy are inversely proportional to the minimum eigenvalue of the predictor sample covariance matrix. Thus, highly correlated predictors deteriorate estimation and prediction performance. Moreover,  Zhao and Yu (2006) proved that the LASSO struggles to differentiate between relevant (i.e., $\{i\in\{1,\dots,n\}:\alpha_{i}^{*}\neq 0\}$) and irrelevant (i.e., $\{i\in\{1,\dots,n\}:\alpha_{i}^{*}=0\}$) predictors when they are closely correlated, leading to false positives. Thus, highly correlated predictors may also deteriorate feature selection performance. The irrepresentable condition addresses this issue ensuring both estimation and feature selection consistency through bounds on the sample correlations between relevant and irrelevant predictors (Zhao and Yu, 2006, see also  Bühlmann and van de Geer, 2011). Nevertheless, orthogonality or weak correlation seldom hold in the context of economic and financial data. For instance, decades of literature provide evidence for co-movements of macroeconomic variables (Forni et al., 2000, 2005; Stock and Watson, 2002a, b). Special methods have been proposed to mitigate the negative effects of these linear associations, such as Factor-Adjusted Regularized Model Selection (FarmSelect) (Fan et al., 2020), which applies the LASSO to the idiosyncratic components of economic variables, obtained by filtering the variables through a factor model. Although approaches such as FarmSelect can be very effective in addressing multicollinearity, strong spurious correlations can emerge at the sample level and affect the LASSO even when predictors are orthogonal or weakly correlated at the population level. Sample-level spurious correlations can be particularly prominent in regressions with many predictors, especially if the sample sizes are relatively small, and the problem can be yet more serious for time series data, where independent replication can be further hindered by serial correlations. This is exactly the focus of this article; in the next section, we introduce a theoretical result linking serial correlations within time series to the sample correlations between them.
@@ 2.1 Serial and Sample Correlations for Time Series
Consider a first order $n$-variate autoregressive process $\mathbf{x}_{t}=\boldsymbol{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$, $t=1,\dots,T$, where $\boldsymbol{\phi}$ is the $n\times n$ diagonal matrix with $\text{diag}(\boldsymbol{\phi})=(\phi_{1},\dots,\phi_{n})$, $|\phi_{i}|<1$ for each $i=1,\dots,n$, and $\mathbf{u}_{t}\sim N\mathopen{}\mathclose{{}\left(\boldsymbol{0}_{n},\boldsymbol{I}_{n}}\right)$. Here $\mathbf{x}_{0}\sim$ $N\mathopen{}\mathclose{{}\left(\boldsymbol{0}_{n},\mathbf{C}_{x}}\right)$ and $\mathbf{x}_{t}\sim$ $N\mathopen{}\mathclose{{}\left(\boldsymbol{0}_{n},\mathbf{C}_{x}}\right)$ with $\mathopen{}\mathclose{{}\left(\mathbf{C}_{x}}\right)_{ii}=\frac{1}{1-\phi_{i}^{2}}$, and $\mathopen{}\mathclose{{}\left(\mathbf{C}_{x}}\right)_{ij}={c}_{ij}^{x}=0$, for $i\neq j$. Let $\widehat{\mathbf{C}}_{x}=\frac{1}{T}\mathbf{X}\mathbf{X}^{\prime}$ be the sample covariance, or equivalently, correlation matrix – with generic off-diagonal element $\widehat{c}_{ij}^{x}$ and eigenvalues $\widehat{\psi}_{max}^{x}\geq\ldots\geq\widehat{\psi}_{min}^{x}$. Our next task is to link $\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)$, $\tau\in[0,1)$, to serial correlations. To this end, the following theorem derives the asymptotic probability density of the sample correlation.
Theorem 1:
Let $\{\mathbf{x}_{t}\}$ be a stationary $n$-variate Gaussian AR(1) process with autoregressive residuals $\mathbf{u}_{t}\sim N\mathopen{}\mathclose{{}\left(\boldsymbol{0}_{n},\boldsymbol{I}_{n}}\right)$. Let $\ddot{\phi}=\phi_{i}\phi_{j}$, where $\phi_{i}$ and $\phi_{j}$ are the autoregressive coefficients of the $i$-th and $j$-th processes, respectively. As $T\rightarrow\infty$, the density of $\widehat{c}_{ij}^{x}$ is
@@
where $T_{v}=\mathopen{}\mathclose{{}\left\lceil\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{i}\phi_{j})^{2}}}\right\rfloor$, $\xi_{v}=3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(1+2\phi_{j}^{2t})$, and $k_{v}=\frac{T_{v}}{\xi_{v}}$.
Proof: see Supplement B.1.
[…]
Figure 1: Monte Carlo generated densities $d(r)$ (top) and asymptotic densities $\mathcal{D}(r)$ (bottom) of the sample correlation $\widehat{c}_{ij}^{x}$ for various values of $T$ and $\phi$.

### v3

@@ 1 Introduction
The LASSO (tibshirani96) is perhaps the most commonly employed approach to handle regressions with a large number of predictors. From a theoretical standpoint, its effectiveness in terms of estimation, prediction, and feature selection is contingent upon either orthogonality or reasonably weak correlation among predictors (see zhaoyu2006; Bickel2009; Negahban2012; Hastie2015). This hinders the use of the LASSO for the analysis of economic time series data, which are notoriously characterized by intrinsic multicollinearity; that is, by predictor correlations at the population level (Lippi2000; SeW2002a; demol2008; Medeiros2012). A common procedure to address this issue is to model multicollinearity and remove it, as proposed, e.g., by Fan2020, who filter time series using common factors and then apply the LASSO to the filtered residuals. However, mitigating or even eliminating multicollinearity is not the end of the story, as effectiveness of the LASSO can also be affected by spurious correlations. These occur when predictors are orthogonal or weakly correlated at the population level, but a lack of sufficient independent replication (lack of degrees of freedom) introduces correlations at the sample level, potentially leading to false scientific discoveries and incorrect statistical inferences (Fan16). This issue has been broadly explored in ultra-high dimensional settings, where the number of predictors can vastly exceed the available sample size (Fan13). We argue that in time series data, a shortage of independent replication can be due not only to a shortage of available observations but also to serial correlation.
@@
Second, we propose an approach that, using a filter similar to that proposed by Fan2020, rescues the performance of the LASSO in the presence of serially correlated predictors. Our approach, which we name ARMAr-LASSO (ARMA residuals LASSO), relies upon a working model where, instead of the observed predictor time series, we use as regressors the residuals of ARMA processes fitted on such series, augmented with lags of the dependent variable. We motivate our choice of working model and provide some asymptotic arguments concerning limiting distribution and feature selection consistency. Next, we employ the mixingale and near-epoch dependence framework (davidson1994; ADAMEK2023) to prove oracle inequalities for the estimation and forecast error bounds of the ARMAr-LASSO, while simultaneously addressing the issue of estimating ARMA residuals. To complete the analysis, we use simulations to validate and generalize theoretical results. Furthermore, we apply our methodology to a high-dimensional dataset for forecasting the consumer price index in the Euro Area. Simulations and empirical exercises demonstrate that the ARMAr-LASSO produces more parsimonious models, better coefficient estimates, and more accurate forecasts than LASSO-based benchmarks. Notably, both theoretical and numerical results concerning our approach hold even in the presence of factor-induced multicollinearity, provided that the idiosyncratic components are orthogonal or weakly correlated processes exhibiting serial correlation.
On the serial correlation front, most of the theoretical econometric literature has focused on its impact in the error terms, particularly regarding post-LASSO inference (see, e.g., Chernozhukov2021; chronopoulos2023; Babii2024; ADAMEK2023). The present study shows that serial correlation in the predictors deserves similar attention, as it can adversely affect both the estimation and forecast accuracy of the LASSO. Our work complements the vast literature on error bounds for LASSO-based methods in time series analysis, which addresses estimation and forecast consistency in scenarios with autocorrelated errors and autoregressive processes (see, e.g., nardi2011; Uematsu2019). Such scenarios are ubiquitous, e.g., they are easily found in US and Euro Area monthly macroeconomic data (see McCracken2016 and Ale2021). Moreover, our methodology is consistent with the existing literature on pre-whitening filters, which aim to mitigate autocorrelation and multicollinearity by applying LASSO or related methods to filtered residuals (see, e.g., Robinson88; Belloni2013; Hansen2019; Fan2020). In particular, the Generalized Least Squares LASSO (GLS-LASSO; chronopoulos2023) and the AutoRegressive Distributed Lag LASSO (ARDL-LASSO; medeiros2017ADA) provide two natural benchmarks for our method. GLS-LASSO improves efficiency by filtering both the dependent variable and the predictors using autoregressive coefficients estimated from residuals of a preliminary LASSO fit, while ARDL-LASSO addresses serial correlation by including lags of both the predictors and the dependent variable. In this paper, we argue that ARMAr-LASSO is preferable to these benchmarks, as it more effectively removes serial correlation in the predictors and thereby enhances both estimation and forecasting performance.
@@
The remainder of the article is organized as follows. Section 2 introduces the problem setup and our results concerning the link between serial correlations and spurious correlations. Section 3 introduces the ARMAr-LASSO and explores its theoretical properties. Section 4 presents simulations and real data analyses to evaluate the proposed methodology. Section 5 provides some final remarks. Appendix A and B contain the proofs of theoretical results and technical details. The Supplement encompasses additional studies and simulations excluded from the main manuscript.
[…]
@@ 2 Problem Setup
where $\mathbf{x}_{t}=(x_{1,t},\ldots,x_{n,t})^{\prime}$ is a $n\times 1$ vector of predictors, $\boldsymbol{\alpha}^{*}$ is a $n\times 1$ unknown $s$-sparse vector of regression coefficients, i.e. $\boldsymbol{\alpha}^{*}||_{0}=s<n$, and $\varepsilon_{t}$ is an error term. We impose the following assumptions on the processes $\mathopen{}\mathclose{{\left\{\mathbf{x}_{t}}}\right\}$ and $\mathopen{}\mathclose{{\left\{\varepsilon_{t}}}\right\}$.
@@
(b) The innovation processes $u_{i,t}\sim i.i.d.(0,\sigma_{i}^{2})$, $\omega_{t}\sim i.i.d.(0,\sigma_{\omega}^{2})$, where $u_{i,t}\perp u_{j,t-l}$ for any $i\neq j,\ t$ and $l\neq 0$; and $u_{i,t-l}\perp\omega_{t}$ for any $i,t$ and $l$.
@@
There are several approaches to estimate a sparse $\boldsymbol{\alpha}^{*}$ (zhang2012; James2013); here we focus on the LASSO estimator (tibshirani96) given by $\widehat{\boldsymbol{\alpha}}=\underset{\boldsymbol{\alpha}\in\mathbb{R}^{n}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{\left\{\frac{1}{2T}||\mathbf{y}-\mathbf{X}^{\prime}\boldsymbol{\alpha}||_{2}^{2}+\ddot{\lambda}||\boldsymbol{\alpha}||_{1}}}\right\},$ where $\mathbf{y}=(y_{1},\dots,y_{T})^{\prime}$ is the $T\times 1$ response vector, $\mathbf{X}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$ is the $n\times T$ design matrix, and $\ddot{\lambda}>0$ is the weight of the $\ell_{1}$ penalty and must be “tuned” to guarantee that regression coefficient estimates are effectively shrunk to zero – thus ensuring predictor, or feature, selection.
@@
However, linear associations among predictors are well known to affect LASSO performance. Bickel2009; Buhlmann2011 and Negahban2012 have shown that the LASSO estimation and prediction accuracy are inversely proportional to the minimum eigenvalue of the predictor sample covariance matrix. Thus, highly correlated predictors deteriorate estimation and prediction performance. Moreover, zhaoyu2006 proved that the LASSO struggles to differentiate between relevant (i.e., $\{i\in\{1,\dots,n\}:\alpha_{i}^{*}\neq 0\}$) and irrelevant (i.e., $\{i\in\{1,\dots,n\}:\alpha_{i}^{*}=0\}$) predictors when they are closely correlated, leading to false positives. Thus, highly correlated predictors may also deteriorate feature selection performance. The irrepresentable condition addresses this issue ensuring both estimation and feature selection consistency through bounds on the sample correlations between relevant and irrelevant predictors (zhaoyu2006, see also Buhlmann2011). Nevertheless, orthogonality or weak correlation seldom hold in the context of economic and financial data. For instance, decades of literature provide evidence for co-movements of macroeconomic variables (Lippi2000; Lippi2005; SeW2002a; SeW2002b). Special methods have been proposed to mitigate the negative effects of these linear associations, such as Factor-Adjusted Regularized Model Selection (FarmSelect) (Fan2020), which applies the LASSO to the idiosyncratic components of economic variables, obtained by filtering the variables through a factor model. Although approaches such as FarmSelect can be very effective in addressing multicollinearity, strong spurious correlations can emerge at the sample level and affect the LASSO even when predictors are orthogonal or weakly correlated at the population level. Sample-level spurious correlations can be particularly prominent in regressions with many predictors, especially if the sample sizes are relatively small, and the problem can be yet more serious for time series data, where independent replication can be further hindered by serial correlations (Bartlett35; Mcgregor1965). This is exactly the focus of this article; in the next section, we introduce a theoretical result linking serial correlations within time series to the sample correlations between them.
@@ 2.1 Serial and Sample Correlations for Time Series
Consider a first order $n$-variate autoregressive process $\mathbf{x}_{t}=\boldsymbol{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$, $t=1,\dots,T$, where $\boldsymbol{\phi}$ is the $n\times n$ diagonal matrix with $\text{diag}(\boldsymbol{\phi})=(\phi_{1},\dots,\phi_{n})$, $|\phi_{i}|<1$ for each $i=1,\dots,n$, and $\mathbf{u}_{t}\sim N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n},\boldsymbol{I}_{n}}}\right)$. Here $\mathbf{x}_{0}\sim$ $N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n},\mathbf{C}_{x}}}\right)$ and $\mathbf{x}_{t}\sim$ $N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n},\mathbf{C}_{x}}}\right)$ with $\mathopen{}\mathclose{{\left(\mathbf{C}_{x}}}\right)_{ii}=\frac{1}{1-\phi_{i}^{2}}$, and $\mathopen{}\mathclose{{\left(\mathbf{C}_{x}}}\right)_{ij}={c}_{ij}^{x}=0$, for $i\neq j$. Let $\widehat{\mathbf{C}}_{x}=\frac{1}{T}\mathbf{X}\mathbf{X}^{\prime}$ be the sample covariance, or equivalently, correlation matrix – with generic off-diagonal element $\widehat{c}_{ij}^{x}$ and eigenvalues $\widehat{\psi}_{max}^{x}\geq\ldots\geq\widehat{\psi}_{min}^{x}$. Our next task is to link $\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)$, $\tau\in[0,1)$, to serial correlations. To this end, the following proposition provides an approximation to the probability density of the sample correlation, yielding a formulation that is simpler than that of Mcgregor1965 and builds upon the results of Anderson03 for i.i.d. random variables.
Proposition 1:
Let $\{\mathbf{x}_{t}\}$ be a stationary $n$-variate Gaussian AR(1) process with autoregressive residuals $\mathbf{u}_{t}\sim N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n},\boldsymbol{I}_{n}}}\right)$. Let $\ddot{\phi}=\phi_{i}\phi_{j}$, where $\phi_{i}$ and $\phi_{j}$ are the autoregressive coefficients of the $i$-th and $j$-th processes, respectively. For some $\nu\in\mathbb{Z}^{+}$, and for all sample sizes $T\geq\mathopen{}\mathclose{{\left\lfloor\nu\mathopen{}\mathclose{{\left(\frac{1+\ddot{\phi}}{1-\ddot{\phi}}}}\right)}}\right\rceil$, the density of $\widehat{c}_{ij}^{x}$ is approximated by
@@
where $T_{v}=\mathopen{}\mathclose{{\left\lfloor\frac{(T-1)(1-\ddot{\phi})^{2}-(1-\ddot{\phi}^{2})}{(1-\ddot{\phi})^{2}}}}\right\rceil$, $\xi_{v}=3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(T_{v}-t)(1+2\phi_{j}^{2t})$, and $k_{v}=\frac{T_{v}}{\xi_{v}}$.
Remark 1:
Proposition 1 establishes a lower bound on the sample size at which $\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)\approx\int_{-1}^{-\tau}\mathcal{D}(r)dr+\int_{\tau}^{1}\mathcal{D}(r)dr$. The bound depends on two quantities: $\nu$, which represents the degrees of freedom under independence (henceforth effective degrees of freedom); and the factor $\mathopen{}\mathclose{{\left(\frac{1+\ddot{\phi}}{1-\ddot{\phi}}}}\right)$, which corresponds to the Bartlett correction for AR(1) processes. The dependence on $\ddot{\phi}$ indicates that, due to serial correlation, the effective degrees of freedom—and thus the effective sample size—is smaller than the nominal sample size (see, e.g., Bartlett35). Note that, for any $|\ddot{\phi}|<1$, $\mathcal{D}(r)$ converges to the Normal distribution as $T\to\infty$.
Figure 1 reports the densities of $\widehat{c}_{ij}^{x}$, indicated as $d(r)$, obtained through 5000 Monte Carlo simulations considering $\phi_{i}=\phi_{j}=0.3,0.6,0.9,0.95$ and $T=50,100,250$. For any $T$ value, an increase in $\ddot{\phi}=\phi_{i}\phi_{j}=\phi^{2}$ results in a density with thicker tails, and thus in a higher $\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)$. The results confirm that serial correlation increases the probability of spurious correlations. This, in turn, leads to a higher probability of a small minimum eigenvalue (because $\Pr(\widehat{\psi}_{min}^{x}\leq 1-\tau)\geq\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)$; see Appendix B), and to a higher chance of breaking the irrepresentable condition if, say, one of the processes is relevant for the response and the other is not ($\alpha_{i}^{*}\neq 0$ and $\alpha_{j}^{*}=0$, or vice versa). Note that this happens when $Sign(\phi_{i})=Sign(\phi_{j})$. In contrast, when $Sign(\phi_{i})\neq Sign(\phi_{j})$, an increase in $|\ddot{\phi}|$ results in a density more concentrated around the origin. In Supplement B, we report a detailed analysis of the results in Figure 1. Furthermore, we investigate the impact of $Sign(\ddot{\phi})$, and more scenarios with correlated, non-Gaussian, and/or ARMA processes, through multiple simulation experiments.
Figure 2 compares $d(r)$ (blue histograms and dashed lines) with $\mathcal{D}(r)$ (red line) considering $\nu=20$ and $\phi_{i}=\phi_{j}=0.3,0.6,0.9$. We observe that $d(r)$ is well approximated by $\mathcal{D}(r)$, indicating that Proposition 1 allows us to explicitly link the probability of sizeable spurious correlations to serial correlations. Note that in the proof of Lemma 1 we provide a theoretical justification for considering $\nu=20$.
Figure 1: Monte Carlo densities $d(r)$ of $\widehat{c}_{ij}^{x}$ for different values of $T$ and $\phi$.
Figure 2: Monte Carlo densities $d(r)$ (blue histograms and dashed lines) and $\mathcal{D}(r)$ (red lines) for $\nu=20$ and different values of $\phi$. The $p$-values correspond to the Shapiro test for Gaussianity.
@@
2.2 LASSO Oracle Inequalities for Orthogonal AR(1) Gaussian Processes
In this section, we establish a connection between LASSO performance and serial correlation, building upon the results derived in Section 2.1. For consistency of exposition, we assume that both the predictors and the error terms are independent Gaussian AR(1) processes.
Assume that each row of $\mathbf{X}$ is standardized to have mean 0 and variance 1. Let $\widehat{\mathbf{C}}_{x}\hskip 2.84544pt\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}\mathbf{C}_{x}$, where $\mathbf{C}_{x}$ is a non-negative definite matrix. In settings where $n$ may be larger than $T$, we usually make the following assumption.
Assumption 2:
[…]
Remark 2:
Since the compatibility constant $\gamma_{x}^{2}$ is directly linked to the smallest eigenvalue of the sample covariance matrix, an increase in predictors’ sample correlations drives the minimum eigenvalue toward zero (see Figure 3 and Appendix B), thereby weakening the RE condition.
Proposition 2:
[…]
Corollary 1:
As a consequence of Proposition 2 we have that: ($a$) $\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{X}^{\prime}(\widehat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{*})}}\right|}}\right|_{2}^{2}=O_{P}\mathopen{}\mathclose{{\left(s{\sigma_{x\varepsilon}^{2}\frac{log(n)}{T}}}}\right)$; and ($b$) $\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{*}}}\right|}}\right|_{1}=O_{P}\mathopen{}\mathclose{{\left(s\sigma_{x\varepsilon}\sqrt{\frac{log(n)}{T}}}}\right)$, where $\sigma_{x\varepsilon}^{2}=\frac{1-\phi^{2}\phi_{\varepsilon}^{2}}{(1-\phi_{\varepsilon}^{2})(1-\phi\phi_{\varepsilon})^{2}}$.
Proposition 2 and Corollary 1 show that oracle inequalities and convergence rates for the LASSO critically depend on $\sigma_{x\varepsilon}^{2}$ and $\gamma_{x}^{2}$. In turn, the results in Section 2.1 show that these key quantities are affected by serial correlation. Hence, as $\phi$ and $\phi_{\varepsilon}$ increase, the bounds in Proposition 2 become larger and the convergence rates in Corollary 1 become slower. Moreover, an increase in $\phi$ leads to an increase in sample correlations (see Proposition 1), which in turn leads to a reduction of $\gamma_{x}^{2}$ (see Remark 2), further amplifying the bounds in Proposition 2.